shasha12:

Which equation correctly applies the distributive property? ​ −2.5⋅(4⋅3.67)=(−2.5⋅4)⋅3.67 ​ ​ 2.4⋅(−3.4)⋅(−1.25)=−3.4⋅2.4⋅(−1.25) ​ ​ (42⋅3.5)+(42⋅1.3)+(42⋅5.2)=42(3.5+1.3+5.2) ​ ​ −3⋅(6.48)=(−3⋅6)+(−3⋅0.4)+(−3⋅0.08) ​

563blackghost:

Distributive Property follows as: \(\Large\bf{a(b+c) = ab + ac}\) Which do you believe follows this format?

shasha12:

the first one?

563blackghost:

Not quite. The first one only distributes once and multiplies instead of adding. \(\large\bf{-2.5 (4 \times 3.67)= (-2.5 \times 4) \color{red}{+} (\color{red}{-2.5 \times~}3.67)}\)

shasha12:

so that leaves the last two

563blackghost:

yes. so is it C or D?

shasha12:

I think C

563blackghost:

Correct!

shasha12:

thank you

563blackghost:

no problem cx